A study of 60 deliveries found a mean time of 24 min, with a standard deviation of 9. a. Calculate a 90% confidence interval for the mean time for a delivery b. Calculate a 90% confidence interval for a single delivery
Answer #1

Solution, Given that, X = 24 S = 9 h - 60 c = 90% = 0.90 At 90% confidence interval t is = 1-C = -0.90 0.1 a = 0.05 Ia 2 dif.